• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name jwexler

    Log entry time 15:53:29 on June17,2010

    Entry number 327085

    keyword=Shift summary - Day 6/16

    Shift Summary:

    08:00 - Beam tuning at start of shift. On empty target, cooler on. Plan is to do short beam recovery.

    08:07 - MCC sending CW.

    08:30 - RC says techs want to go into the hall for dipole-related work. Called for controlled access.

    09:35 - Right Q1 tripped. Power cycling it remotely looks to be working.

    12:00 - Decide to warm up target. JP begins process.

    13:20 - Heidi came in, says that the magnets could go down when putting the target back on the cold lines. We ramped down the magnets. The regulation shouldn't be turned back on until after the target is cold.

    15:00 - Ed going into the hall to turn on the pumps

    15:45 - Look at target camera, see the target looks like the ice is gone.

    Runs:
    Pedestal - 4627-4634